# WebVoyager [Uncaptioned image]: Building an End-to-End Web Agent with Large Multimodal Models

### 25 Jan 2024 v1

@@ Abstract
The advancement of large language models (LLMs) leads to a new era marked by the development of autonomous applications in the real world, which drives innovation in the creation of advanced web-based agents. Existing web agents typically only handle one input modality and are evaluated only in simplified web simulators or static web snapshots, greatly limiting their applicability in real-world scenarios. To bridge this gap, we introduce WebVoyager, an innovative Large Multimodal Model (LMM) powered web agent that can complete user instructions end-to-end by interacting with real-world websites. Moreover, we propose a new evaluation protocol for web agents to address the challenges of automatic evaluation of open-ended web agent tasks, leveraging the robust multimodal comprehension capabilities of GPT-4V. We create a new benchmark by gathering real-world tasks from 15 widely used websites to evaluate our agents. We show that WebVoyager achieves a 55.7% task success rate, significantly surpassing the performance of both GPT-4 (All Tools) and the WebVoyager (text-only) setups, underscoring the exceptional capability of WebVoyager in practical applications. We found that our proposed automatic evaluation achieves 85.3% agreement with human judgment, paving the way for further development of web agents in a real-world setting. [^1] Our code and data will be released at https://github.com/MinorJerry/WebVoyager
@@ 1 Introduction
However, existing approaches overlook a critical functionality of browsing: rendering HTML into visual webpages. Particularly, vision capability is crucial for utilizing tools like web browsers, as rendered web pages are inherently designed with user experience (UX), emphasizing intuitive information and structured presentation. This design principle of rendering makes visual analysis more effective than mere HTML representation. At present, large multimodal models (LMMs), particularly GPT-4V(ision) (OpenAI, 2023) and Gemini (Team et al., 2023), demonstrate a remarkable ability to integrate intricate visual cues with textual information. Existing studies such as Pix2Struct (Lee et al., 2023) and WebArena (Zhou et al., 2023), have initiated explorations into using screenshots as inputs for decision-making in web navigation, yet these are preliminary and do not represent a deep exploration. Therefore, building multimodal web agents to leverage the environment rendered by browsers through screenshots, thus mimicking human web browsing behavior, is now a viable approach to enhance web navigation efficiency.
@@
Figure 1: The overall workflow of WebVoyager. Humans assign web tasks to WebVoyager, which automatically browses the web online. At each step, WebVoyager selects actions based on screenshots and text (the ‘type’ of the web element and its contents). Once the task is completed, the answers will be returned to the user. For example, a user queries: "Find the cost of a 2-year protection for PS4 on Amazon." Agent interacts with Amazon online, locates the PS4, identifies the 2-year protection price, and returns "\$30.99" to the user.
We introduce WebVoyager, a multimodal web agent designed to handle web tasks online in an end-to-end manner, which denotes managing the process from start to finish autonomously without intermediate human intervention. We construct an online environment using Selenium for WebVoyager, feeding it with screenshots and textual content in interactive web elements. Inspired by Set-of-Mark Prompting (Yang et al., 2023a), we mark interactive web elements on screenshots (see Figure 2) to facilitate decision-making for WebVoyager. As a pioneer in combining vision and text information during web navigation, we advocate that autonomous end-to-end task completion, multimodal capabilities and online navigation constitute the essential trajectory toward the genuine intelligence of web agents.
@@
Another challenge arises when it comes to evaluating an end-to-end web agent with online navigation. Existing benchmarks, such as Mind2Web (Deng et al., 2023), primarily focus on stepwise and offline evaluation, where agents follow predefined “golden” trajectory for action selection. This approach, however, may not fully account for the variety of viable strategies to accomplish a task, as it only reflects one possible plan. This limitation could lead to a biased evaluation and difficulties in fairly comparing different methods. To more accurately gauge the capabilities of web agents in end-to-end task completion, we save screenshots throughout the online navigation process, and then use GPT-4V to evaluate these trajectories and the final results automatically. Human evaluations are also conducted to verify the results and confirm the reliability of GPT-4V as the evaluator.
@@
We conduct evaluations on a collected dataset, which is semi-automatically generated using a self-instruct (Wang et al., 2022) method, comprising 300 web tasks from 15 commonly accessed websites. Additionally, we extract 90 web-related tasks of level 1 and level 2 from the GAIA (Mialon et al., 2023) to enrich our evaluation. We compare our WebVoyager with 1) GPT-4 (All Tools) [^2] GPT-4 (All Tools) is an integrated tool-based agent released by OpenAI in Oct. 2023. See https://chat.openai.com/, and 2) WebVoyager in a text-only setting, employing the accessibility tree proposed in WebArena (Zhou et al., 2023) to describe web pages. The results show that WebVoyager achieves a Task Success Rate of 55.7%, significantly outperforming GPT-4 (All Tools) with a rate of 32.7% and the text-only setting with a rate of 39.0%, demonstrating the effectiveness of our method. Furthermore, we report the consistency between human-human and human-GPT4V to ensure credibility. Our main contributions are as follows:
  • We employ a multimodal web agent that integrates textual and visual information to address web tasks end-to-end and introduce a generalist planning approach for navigation.
  • We build an online web browsing environment, offering a variety of tasks centered on widely used websites and introducing a method for expanding these tasks.
  • We conduct manual evaluations of navigation trajectories and propose an automated evaluation protocol using GPT-4V. We present a comprehensive analysis of the evaluation results and show that GPT-4V can serve as a reliable evaluator for online agents.
@@ 2.1 Web Agents
Along with these new benchmarks, numerous efforts have been made to build autonomous web agents. WebGPT (Nakano et al., 2021) constructs a text-based web browsing environment and fine-tunes GPT-3 as a web agent. WebAgent (Gur et al., 2023) pretrains a T5 model to extract HTML snippets and leverages Flan-U-PaLM (Chowdhery et al., 2023) to generate Python code to interact with the environment. Besides fine-tuning, another line of work tries to build web agents by prompting LLMs Yao et al. (2022b); Shinn et al. (2023); Ma et al. (2023). Multimodal web agents that integrate visual signals have also been explored, WebGUM (Furuta et al., 2023) combines T5 (Raffel et al., 2020) with a Vision Transformer (ViT) to navigate using both screenshots and HTML text. PIX2ACT (Shaw et al., 2023) instead solely relies on web screenshots as inputs to predict agent actions. Different from previous works that only consider a single modality or simplified web environments, in this work, we build a multimodal agent that can complete tasks on real-world websites. Concurrent to our work, SEEACT (Zheng et al., 2024) implements a multimodal web agent leveraging GPT-4V, but they mainly focus on offline navigation where the agent observes the static website snapshot to make predictions. We show that our agent can complete user tasks and get up-to-date answers by interacting with a web browser.
2.2 Large Multimodal Models
In recent years, significant strides have been made in unifying image and text representations within a single multimodal model through joint training with image and text (Li et al., 2019; Dosovitskiy et al., 2020; Wang et al., 2021; Dai et al., 2022; Aghajanyan et al., 2022). Large Multimodal Models (LMMs), following in the footsteps of Large Language Models (Brown et al., 2020; Chen et al., 2021; Chowdhery et al., 2023), attain the capability of instruction following (Ouyang et al., 2022) and exhibit robust multimodal comprehension. Represented by GPT-4V (OpenAI, 2023) and Gemini (Team et al., 2023), LMMs have demonstrated impressive performance on benchmarks (Goyal et al., 2017; Lu et al., 2022; Zellers et al., 2019; Hessel et al., 2022), establishing a foundation for the construction of multimodal agents in subsequent research.
@@ 3 WebVoyager
We aim to build an agent that can browse the open web autonomously without any human intervention to complete user instructions. Given an instruction, our WebVoyager first instantiates a web browser as the interface to the open web. Similar to humans, WebVoyager takes the visual signals from the web (e.g. screenshots) as the primary source of input at every step, and it considers text extracted from key HTML elements in addition. The agent produces an action based on the inputs at every step, which is then executed in the browser environment. The process continues until the agent decides to stop. Next, we describe the details of WebVoyager, including environment, interaction cycle, observation space, and action space.
@@ 3.1 Browsing Environment
We develop an automated web-browsing environment using Selenium, a powerful tool capable of emulating user browsing behaviors. Our environment provides the same level of website accessibility to the agent as a human user, i.e. a browser window, similar to WebArena Zhou et al. (2023). Unlike WebArena, we do not host any websites locally and allow the agent to explore the open web instead. Unlike locally hosted websites, open web poses unique challenges, e.g. floating ads, pop-up windows, constant updates, etc. [^3] Regarding CAPTCHAs (Completely Automated Public Turing test to tell Computers and Humans Apart) challenges, we believe it is important to respect the rules of these websites and prompt the agent to retrieve information from alternative sources. Still, we opt for online interaction with real websites for our agent as we believe this setting truly reflects the real-world use cases (e.g. the agent needs access to real-time information from the web) and we hope to build the agent that can robustly adapt to these challenges and consistently solve the problem.
@@ 3.2 Interaction Formulation
Formally, we denote the Environment as $\mathcal{E}$, the large Multimodal Model as $\mathcal{M}$, the Observation Space as $\mathcal{O}$, and the Action Space as $\mathcal{A}$. At time step $t$, the model receives the context $c_{t}$ as inputs, which consists of historical actions $a_{t}$ and observations $o_{t}$, defined as: $c_{t}=(o_{1},a_{1},...,o_{t-1},a_{t-1},o_{t},I)$ The the model produces the action $a_{t}$ at time $t$, $a_{t}=\mathcal{M}(c_{t})$, which is then executed in the environment. After execution, the environment sends back the observation at time $t+1$, $o_{t+1}=\mathcal{E}(o_{t},a_{t})$. Then the context will be updated and this interaction process continues until the model generates a terminating action or the maximum step is reached.
Inspired by the paradigm of ReAct Prompting (Yao et al., 2022b), we also prompt our agent to generate a thought process first before generating the action code. Hence $(s_{t},a_{t})$ can be further composed into $\hat{a}_{t}$ where $s_{t}$ and $a_{t}$ represent the natural language thought and action code respectively. Figure 7 in Appendix A presents the System Prompt we designed for the action prediction step. Also, it’s worth noting that excessive screenshots of web pages from longer episodes may confuse the agent. Therefore, we perform context clipping to remove outdated web page information and only keep the most recent 3 screenshots in the inputs and we keep the entire history of thoughts and actions to better guide the agent.
@@ 3.3 Observation Space
Similar to how humans browse the web, our agent also takes the visual information from the web (screenshots) as the primary source of input. Using screenshots allows us avoid the burden of processing HTML DOM tree or accessibility tree to portray the overall structure of webpages, which can lead to overly verbose texts and impact the decision-making process of the agent. Inspired by Set-of-Mark Prompting (Yang et al., 2023a), we overlay bounding boxes of the interactive elements on the websites to better guide the agent’s action prediction. Unlike Yang et al. (2023a), we do not need any object detection module (Zou et al., 2023). Instead, we utilize GPT-4-ACT [^4] https://github.com/ddupont808/GPT-4V-Act, a Javascript tool to extracts the interactive elements based on web element types and then overlays bounding boxes with numerical labels on the respective regions of the elements. As illustrated in Figure 2, the nature of webpages allows us to precisely locate and outline each interactive element using this tool. The numerical labels assigned to each element are also essential for the model to identify the elements requiring interaction, thereby facilitating accurate action determination. We empirically choose black as the color for the borders and the background of the labels to enhance clarity. We observe that using a single black color yields higher success rates in planning compared to using multiple colors. Additionally, we also provide the agent some auxiliary text as inputs, including the textual content embedded within the interactive element, the type of the element, and possibly some comment text in the aria-label attribute. To simplify the observation, we have disabled multiple tabs, i.e. all interactions occur within the current tab instead of opening new ones.
At every step, the agent receives current screenshot and auxiliary text along with history as inputs, as discussed in (§3.2). In case the agent’s action raised an exception during execution, we additionally incorporated the error messages in the prompt and ask the model to regenerate the response. Note that each of the error correction attempts also consume one step from the total exploration budget.

### 29 Feb 2024 v3

@@ Abstract
The rapid advancement of large language models (LLMs) has led to a new era marked by the development of autonomous applications in real-world scenarios, which drives innovation in creating advanced web agents. Existing web agents typically only handle one input modality and are evaluated only in simplified web simulators or static web snapshots, greatly limiting their applicability in real-world scenarios. To bridge this gap, we introduce WebVoyager, an innovative Large Multimodal Model (LMM) powered web agent that can complete user instructions end-to-end by interacting with real-world websites. Moreover, we establish a new benchmark by compiling real-world tasks from 15 popular websites and introduce an automatic evaluation protocol leveraging multimodal understanding abilities of GPT-4V to evaluate open-ended web agents. We show that WebVoyager achieves a 59.1% task success rate on our benchmark, significantly surpassing the performance of both GPT-4 (All Tools) and the WebVoyager (text-only) setups, underscoring the exceptional capability of WebVoyager. The proposed automatic evaluation metric achieves 85.3% agreement with human judgment, indicating its effectiveness in providing reliable and accurate assessments of web agents. [^1] Our code and data will be released at https://github.com/MinorJerry/WebVoyager
@@ 1 Introduction
However, existing approaches overlook a critical functionality of browsing: rendering HTML into visual webpages. Particularly, vision capability is crucial for utilizing tools such as web browsers, as rendered web pages are inherently designed with user experience (UX), emphasizing intuitive information and structured presentation. This design principle of rendering makes visual analysis more effective than mere HTML representation. At present, large multimodal models (LMMs), particularly GPT-4V(ision) (OpenAI, 2023) and Gemini (Team et al., 2023), demonstrate a remarkable ability to integrate intricate visual cues with textual information. Existing studies such as Pix2Struct (Lee et al., 2023) and WebArena (Zhou et al., 2023), have initiated explorations into using screenshots as inputs for decision-making in web navigation, yet these are preliminary and do not represent a deep exploration. Therefore, building multimodal web agents to leverage the environment rendered by browsers through screenshots, thus mimicking human web browsing behavior, is now a viable approach to enhance web navigation abilities.
@@
Figure 1: The overall workflow of WebVoyager. WebVoyager takes web tasks assigned by a human and automatically browses the web online. At each step, WebVoyager selects actions based on screenshots and text (the ‘type’ of the web element and its contents). Once the task is completed, the answers will be returned to the user. For example, a user queries: "Find the cost of a 2-year protection for PS4 on Amazon." Agent interacts with Amazon online, locates the PS4, identifies the 2-year protection price, and returns "\$30.99" to the user.
We introduce WebVoyager (Figure 1), a multimodal web agent designed to autonomously accomplish web tasks online from start to finish, managing the entire process end-to-end without any intermediate human intervention. WebVoyager processes the user query by making observations from screenshots and textual content in interactive web elements, formulates a thought on what action to take (such as clicking, typing, or scrolling, etc.), and then executes that action on the websites. Inspired by Set-of-Mark Prompting (Yang et al., 2023a), we mark interactive web elements on screenshots (see Figure 2) to facilitate decision-making for WebVoyager.
@@
Another challenge is the evaluation of an end-to-end web agent. Existing benchmarks, such as Mind2Web (Deng et al., 2023), primarily focus on stepwise and offline evaluation, where agents follow a predefined “golden” trajectory for action selection. This approach, however, may not fully account for the variety of viable strategies to accomplish a task, as it only reflects one possible plan. This limitation could lead to a biased evaluation and difficulties in fairly comparing different methods. To accurately evaluate the capabilities of web agents in end-to-end task completion, we propose an automated evaluation protocol using GPT-4V. Specifically, we save screenshots throughout the online navigation process and then use GPT-4V to evaluate these trajectories together with the final results automatically. Human evaluations are also conducted to verify the results and the analysis shows that our evaluation protocol achieves 85.3% agreement with human judges, indicating GPT-4V can serve as a reliable evaluator for online agents.
@@
We conduct evaluations on a newly-collected dataset, which is semi-automatically generated using a self-instruct (Wang et al., 2022) method, comprising 643 web tasks from 15 commonly accessed websites. We also evaluate WebVoyager on 90 web-related tasks of level 1 and level 2 from the GAIA (Mialon et al., 2023), and 50 interactive open-web tasks from SeeAct Zheng et al. (2024). We compare our WebVoyager with 1) GPT-4 (All Tools) [^2] GPT-4 (All Tools) is an integrated tool-based agent released by OpenAI in Oct. 2023. See https://chat.openai.com/, and 2) WebVoyager in a text-only setting which employs the textual accessibility tree proposed in WebArena (Zhou et al., 2023) to describe web pages. The results show that WebVoyager achieves a Task Success Rate of 59.1% on our new benchmark, significantly outperforming GPT-4 (All Tools) with a rate of 30.8% and the text-only setting with a rate of 40.1%, demonstrating the effectiveness of our method. Our research demonstrates the effectiveness of the WebVoyager method for web tasks, offering insights into the development of more intelligent and efficient web automation solutions.
@@ 2 Related Work
Along with these new benchmarks, numerous efforts have been made to build autonomous web agents. WebGPT (Nakano et al., 2021) constructs a text-based web browsing environment and fine-tunes GPT-3 as a web agent. WebAgent (Gur et al., 2023) pretrains a T5 model to extract HTML snippets and leverages Flan-U-PaLM (Chowdhery et al., 2023) to generate Python code to interact with the environment. Besides fine-tuning, another line of work tries to build web agents by prompting LLMs Yao et al. (2022b); Shinn et al. (2023); Ma et al. (2023). Multimodal web agents that integrate visual signals have also been explored, WebGUM (Furuta et al., 2023) combines T5 (Raffel et al., 2020) with a Vision Transformer (ViT) to navigate using both screenshots and HTML text. PIX2ACT (Shaw et al., 2023) instead solely relies on web screenshots as inputs to predict agent actions. Unlike previous works that only consider a single modality or simplified web environments, we build a multimodal agent that can complete tasks on real-world websites in this work. Concurrently with our work, SeeAct (Zheng et al., 2024) also leverages Large Multimodal Models (LMMs) for integrated visual understanding and actions on websites. However, the best SeeAct agent still relies on a finetuned cross-encoder model to select candidate elements for interaction. In contrast, WebVoyager do not require any additional modules.
@@ 3 WebVoyager
We aim to build an agent that can browse the open web autonomously without human intervention to complete user instructions. Given an instruction, our WebVoyager first instantiates a web browser and then performs actions with visual (i.e., screenshots) and textual (i.e., HTML elements) signals from the web. The agent produces an action based on the inputs at every step, which is then executed in the browser environment. The process continues until the agent decides to stop. The details of WebVoyager, including environment, interaction cycle, observation space, and action space, are as follows.
@@ 3.1 Browsing Environment
We develop an automated web-browsing environment using Selenium [^3] https://www.selenium.dev/. Unlike WebArena Zhou et al. (2023), we do not host any websites locally and allow the agent to explore the open web instead, which poses unique challenges such as floating ads, pop-up windows, constant updates, etc. [^4] Regarding CAPTCHAs (Completely Automated Public Turing test to tell Computers and Humans Apart) challenges, we believe it is important to respect the rules of these websites and prompt the agent to retrieve information from alternative sources. Still, we opt for online interaction with real websites as we believe that this setting truly reflects the real-world use cases (e.g., the agent needs access to real-time information from the web), and a successful web agent should be able to adapt to these challenges and consistently solve the problem robustly.
@@ 3.2 Interaction Formulation
Formally, we denote the Environment as $\mathcal{E}$, the large Multimodal Model as $\mathcal{M}$, the Observation Space as $\mathcal{O}$, and the Action Space as $\mathcal{A}$. At time step $t$, the model receives the context $c_{t}$ as inputs, which consist of historical actions $a_{i}$ and observations $o_{i}$, defined as: $c_{t}=(o_{1},a_{1},...,o_{t-1},a_{t-1},o_{t},I)$ The the model produces the action $a_{t}$ at time $t$, $a_{t}=\mathcal{M}(c_{t})$, which is then executed in the environment. After execution, the environment sends back the observation at time $t+1$, $o_{t+1}=\mathcal{E}(o_{t},a_{t})$. Then the context will be updated and this interaction process continues until the model generates a terminating action or the maximum step is reached.
Inspired by the paradigm of ReAct Prompting (Yao et al., 2022b), we also prompt our agent to generate a thought process first before generating the action code. Hence $a_{t}$ can be further composed into $(s_{t},\hat{a}_{t})$ where $s_{t}$ and $\hat{a}_{t}$ represent the natural language thought and action code respectively. Figure 7 in Appendix A presents the System Prompt we designed for the action prediction step. Also, it’s worth noting that excessive observations of web pages from longer episodes may confuse the agent. Therefore, we perform context clipping to remove outdated web page information and only keep the three most recent observations in the inputs, and we keep the entire history of thoughts and actions to better guide the agent.
@@ 3.3 Observation Space
Similar to how humans browse the web, our agent also takes the visual information from the web (screenshots) as the primary source of input. Using screenshots allows us to avoid the burden of processing HTML DOM tree or accessibility tree to portray the overall structure of webpages, which can lead to overly verbose texts and impact the decision-making process of the agent. Inspired by Set-of-Mark Prompting (Yang et al., 2023a), we overlay bounding boxes of the interactive elements on the websites to better guide the agent’s action prediction. Unlike Yang et al. (2023a), we do not need any object detection module (Zou et al., 2023). Instead, we utilize GPT-4-ACT [^5] https://github.com/ddupont808/GPT-4V-Act, a Javascript tool to extracts the interactive elements based on web element types and then overlays bounding boxes with numerical labels on the respective regions of the elements.
As illustrated in Figure 2, the nature of webpages allows us to locate and outline each interactive element using this tool precisely. The numerical labels assigned to each element are also essential for the model to identify the elements requiring interaction, thereby facilitating accurate action determination. We empirically choose black color for the borders and the background of the labels to enhance clarity. We observe that using a single black color yields higher success rates than using multiple colors. We also provide the agent with auxiliary text as inputs, including the textual content embedded within the interactive element, the type of the element, and possibly some comment text in the aria-label attribute. To simplify the observation, we have disabled multiple tabs, i.e., all interactions occur within the current tab instead of opening new ones.
At every step, the agent receives the current screenshot, auxiliary text, and history as inputs, as discussed in (§3.2). In case the agent’s action raised an exception during execution, we additionally incorporated the error messages in the prompt and asked the model to regenerate the response. Note that each error correction attempt also consumes one step from the total exploration budget.